Note: Students must show all the work to get full creait. No formula sheet is needed. Each question is 20 points Four point charges “q” are located at the corner of a square of side “a meter. All charges are negative except the one located at the origin. Find: (a) the magnitude and (b) the direction of the net force on the positive charge. 1)
Answer #1

Electric field due to the charge 1
E1 = Kq/ r2
where r is the distance between the charge 1 and the origin.
hence
E1 = kq /a2
Similarly E3   = Kq/a2
And the E2 = Kq / r2
for charge 2,  
r = 21/2a therefore
E2 = Kq/ 2a2
Now We take the component of E2 along Y and Z axis
E2y = E2Cos45
E2z = E2Sin45
Hence the Net electric field in the Z direction
Ez = E1 + E2Sin45 = Kq/a2 + (Kq/ 2a2)Sin45
= 1.353[Kq/a2] -------------(1)
Similarly in the y direction
EY = E2Cos45 + E3
= 1.353[Kq/a2] -------(2)
Hence the magnitude of the Electric field
E = (Ez2 + Ey2)1/2 = 1.914[Kq/a2]
Hence the magnitude of the force = qE
= 1.914[Kq2 /a2]
And the direction of the force would be same as that of field
\theta = tan-1(EZ /EY) = 45
Hence the direction of the force would be 45 from the Y axis in the anticlockwise direction.